lizzie777:

What is the solution to the equation One-half x + 3 = two-thirds x + 1?

ramen:

So the current equation is 1/2x+3=2/3x+1 Do you know how to get the solution of this equation so far?

lizzie777:

i dont know what is 2/3 - 1/2

ramen:

try finding common denominators

ramen:

whats the common denominator of 2/3 and 1/2?

ramen:

or you could just use ur calculator to find decimals. just type into ur calculator (2/3)-(1/2). if you get a decimal its fine

darkknight:

Do it in a fraction. It will make your life easier.

darkknight:

So as ramen said, find the common denominators, then rewrite using the same common denominator. The subtract them

ramen:

anway, the common denominator is 6. I'll let you figure out the numerator

darkknight:

Ramen, don't give direct answers bro. So he said the least common denominator was 6 because the least common denominator is : "the lowest common multiple of the denominators of several fractions." So we need to rewrite the fractions in terms of their least common denominators. 2/3 = 4/6 1/2 = 3/6 So 2/3 - 1/2 is the same as 4/6-3/6. Can you solve for that now? Also remember, when finding solutions to X, you want to isolate x and solve for it. So tell us what you get so far.

ramen:

its not a direct answer. a direct answer is when u give the answer to the whole problem. this was only a part of the problem.

darkknight:

ik, just explain your thought process more in depth so the user can understand. Now lets wait for @lizzie777 's reply

ramen:

yea so what dark said is right. we have to isolate the Xs on one side and the numbers on the other. can you do that?

lizzie777:

yea

darkknight:

Did you get your answer?

ramen:

Whats 4/6x -3/6x = 3-1 ? this is simplified for you.

ramen:

You can't give direct answer unazaaaa

unazaaaa:

oh oops i didnt know im so sorryyyy

ramen:

Its alright but its just against the rules. plz let the lizz figure it out. we can only give her clues

unazaaaa:

okay i'll remember that for next time

Hero:

"oops I didn't know" is no longer an excuse. Read the rules @unazaaaa

unazaaaa:

uhhh lolll im still kinda new here sorry that it was such a problem @Hero